OpenStudy (anonymous):

Find the volume of right circular cone if the lateral surface is to be maximized and is inscribed in a sphere of radius 9 in (Hint: let theta be the central angle subtending a radius of the base of the cone.)

OpenStudy (anonymous):

Let the radius of the cone be r height be h and length be l . \[h = 9 + \sqrt{81 - r ^{2}}\]

OpenStudy (anonymous):

\[l = \sqrt{r ^{2} + h ^{2}}\]

OpenStudy (anonymous):

lateral surface area = pi * r * l write l in terms of r and then differentiate wrt r and put to zero you will get values of r for which lateral surface area is maximum

OpenStudy (anonymous):

why is h = 9 + sqrt rt............?

OpenStudy (anonymous):

@suvesh253